OpenStudy (anonymous):

Which algebraic expression is equivalent to the expression below? 5(8 + 5x) + 11 A. 25x + 51 B. 5x + 51 C. 25x - 51 D. 25x + 29 @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

40+25x+11=51+25x=25x+51

OpenStudy (uri):

5*8+5*5x +11 @rossseeee